OpenStudy (anonymous):

Factor completely 9x^2-6xy+y^2-81

OpenStudy (anonymous):

Steps: 1. factor out 3 in the terms. 2. Group like terms and factor out the common factor

OpenStudy (anonymous):

@Hotchellerae21 almost done :O! ♥♥♥♥♥ only 3 more

OpenStudy (anonymous):

9x² - 6xy + y² - 81 = ( 9x² - 6xy + y² ) - 81 = ( 9x² - 6xy + y² ) - ( 9 )² First bracket = a² - 2ab + b² = ( a - b )² = [ ( 3x )² - 2( 3x )( y ) + ( y )² ] - ( 9 )² = ( 3x - y )² - ( 9 )² Difference of 2 squares = a² - b² = ( a + b )( a - b ) = ( 3x - 9y + 9 )( 3x - 9y - 9 )

OpenStudy (anonymous):

You're amazinggg ♥

OpenStudy (anonymous):

thanks